OpenStudy (anonymous):

Which is the equation of the given line in slope-intercept form? The points that create the line are: (-1,5) and (1,-1) Explain your answer and thanks for the help! A. y=3x+2 B. y=3x-2 C. y=-3x+2 D. y=-1/3x+2

OpenStudy (anonymous):

ello love :D @xDanix need some help?

OpenStudy (sushi121212):

I would think the answer would be C, y=-3x+2 because you first have to find the slope: 5+1 y-y ___ because the formula is ___ -1-1 x-x That would leave you 6/-2 or -3 for the slope. Then, you plug in the points for x and y: y+1=-3(x-1) y+1=-3x+3 -1 -1 ________ y=-3x+2 I hope this helps!

OpenStudy (anonymous):

@sushi121212 It helps a lot! Thank you very much!

OpenStudy (anonymous):

@atreyu6s3x6 I did, but it looks like someone helped me already! But thank you for wanting to help, "love". ;D

OpenStudy (anonymous):

haha it's quite alright lol :) @xDanix